# Alternating Direction Method of Multipliers AR AutoRally

### v1

@@ VI Conclusion
In this paper, we introduce a new sampling-based optimization C++/CUDA library called MPPI-Generic. It contains implementations of MPPI, Tube-MPPI, and RMPPI controllers as well as an API that allows these controllers to be used with multiple dynamics and cost functions. We went through the various ways that researchers could use this library, from using pre-defined dynamics and cost functions to implementing your own sampling-based MPC controller. We also discuss the methods we used to improve the computational performance. We conducted performances comparisons against other widely-available implementations of MPPI over a variety of computer hardware to show the performance benefits our library can provide. Finally, we compared against a sample-efficient form of MPPI to show that with the speed improvements of our library, using more samples is a viable alternative with little hit to computation times. We plan to keep working on the library to add more capabilities and usage improvements such as a Python wrapper in the future.
@@
ARMv8 (v8l) Tegra X1 128 MPPI-Generic 1.024 39±0.091 405 8uncertain1.024390.09140581.024\,39$\pm$ 0.091\,405\,8start_ARG 1.024 39 end_ARG ± start_ARG 0.091 405 8 end_ARG ARMv8 (v8l) Tegra X1 128 autorally 1.396 56±0.613 235uncertain1.396560.6132351.396\,56$\pm$ 0.613\,235start_ARG 1.396 56 end_ARG ± start_ARG 0.613 235 end_ARG
@@
ARMv8 (v8l) Tegra X1 256 MPPI-Generic 1.383 92±0.061 362 6uncertain1.383920.06136261.383\,92$\pm$ 0.061\,362\,6start_ARG 1.383 92 end_ARG ± start_ARG 0.061 362 6 end_ARG ARMv8 (v8l) Tegra X1 256 autorally 1.506 45±0.575 889uncertain1.506450.5758891.506\,45$\pm$ 0.575\,889start_ARG 1.506 45 end_ARG ± start_ARG 0.575 889 end_ARG
@@
ARMv8 (v8l) Tegra X1 512 autorally 1.846 51±0.574 576uncertain1.846510.5745761.846\,51$\pm$ 0.574\,576start_ARG 1.846 51 end_ARG ± start_ARG 0.574 576 end_ARG ARMv8 (v8l) Tegra X1 512 MPPI-Generic 2.122 54±0.119 91uncertain2.122540.119912.122\,54$\pm$ 0.119\,91start_ARG 2.122 54 end_ARG ± start_ARG 0.119 91 end_ARG
@@
ARMv8 (v8l) Tegra X1 1024 autorally 3.244 59±0.607 794uncertain3.244590.6077943.244\,59$\pm$ 0.607\,794start_ARG 3.244 59 end_ARG ± start_ARG 0.607 794 end_ARG ARMv8 (v8l) Tegra X1 1024 MPPI-Generic 4.105 41±1.202 12uncertain4.105411.202124.105\,41$\pm$ 1.202\,12start_ARG 4.105 41 end_ARG ± start_ARG 1.202 12 end_ARG
@@
ARMv8 (v8l) Tegra X1 2048 autorally 5.921±0.577 712uncertain5.9210.5777125.921$\pm$ 0.577\,712start_ARG 5.921 end_ARG ± start_ARG 0.577 712 end_ARG ARMv8 (v8l) Tegra X1 2048 MPPI-Generic 7.296 89±0.398 892uncertain7.296890.3988927.296\,89$\pm$ 0.398\,892start_ARG 7.296 89 end_ARG ± start_ARG 0.398 892 end_ARG
@@
ARMv8 (v8l) Tegra X1 4096 autorally 11.1589±0.581 031uncertain11.15890.58103111.1589$\pm$ 0.581\,031start_ARG 11.1589 end_ARG ± start_ARG 0.581 031 end_ARG ARMv8 (v8l) Tegra X1 4096 MPPI-Generic 14.1242±0.133 369uncertain14.12420.13336914.1242$\pm$ 0.133\,369start_ARG 14.1242 end_ARG ± start_ARG 0.133 369 end_ARG
@@
ARMv8 (v8l) Tegra X1 6144 autorally 16.5458±0.605 249uncertain16.54580.60524916.5458$\pm$ 0.605\,249start_ARG 16.5458 end_ARG ± start_ARG 0.605 249 end_ARG ARMv8 (v8l) Tegra X1 6144 MPPI-Generic 21.0746±0.229 213uncertain21.07460.22921321.0746$\pm$ 0.229\,213start_ARG 21.0746 end_ARG ± start_ARG 0.229 213 end_ARG
@@
ARMv8 (v8l) Tegra X1 8192 autorally 21.7672±0.628 532uncertain21.76720.62853221.7672$\pm$ 0.628\,532start_ARG 21.7672 end_ARG ± start_ARG 0.628 532 end_ARG ARMv8 (v8l) Tegra X1 8192 MPPI-Generic 28.0223±0.260 565uncertain28.02230.26056528.0223$\pm$ 0.260\,565start_ARG 28.0223 end_ARG ± start_ARG 0.260 565 end_ARG
@@
ARMv8 (v8l) Tegra X1 16384 autorally 44.1579±0.947 064uncertain44.15790.94706444.1579$\pm$ 0.947\,064start_ARG 44.1579 end_ARG ± start_ARG 0.947 064 end_ARG ARMv8 (v8l) Tegra X1 16384 MPPI-Generic 56.6497±0.350 869uncertain56.64970.35086956.6497$\pm$ 0.350\,869start_ARG 56.6497 end_ARG ± start_ARG 0.350 869 end_ARG

### v2

@@ VI Conclusion
In this paper, we introduce a new sampling-based optimization C++/CUDA library called MPPI-Generic. It contains implementations of MPPI, Tube-MPPI, and RMPPI controllers as well as an API that allows these controllers to be used with multiple dynamics and cost functions. We went through various ways that researchers could use this library, from using pre-defined dynamics and cost functions to implementing new sampling-based MPC controllers. We discussed the methods used to improve the computational performance and conducted performances comparisons against other widely-available implementations of MPPI over a variety of computer hardware to show the performance benefits our library can provide. Finally, we compared against a sample-efficient form of MPPI to show that with the speed improvements of our library, using more samples is a viable alternative with little hit to computation times. We plan to keep working on the library to add more capabilities and usage improvements such as a Python wrapper in the future.
@@
Jetson Nano Tegra X1 128 autorally 1.396 56±0.613 235uncertain1.396560.6132351.396\,56$\pm$ 0.613\,235start_ARG 1.396 56 end_ARG ± start_ARG 0.613 235 end_ARG Jetson Nano Tegra X1 128 MPPI-Generic 1.024 39±0.091 405 8uncertain1.024390.09140581.024\,39$\pm$ 0.091\,405\,8start_ARG 1.024 39 end_ARG ± start_ARG 0.091 405 8 end_ARG
@@
Jetson Nano Tegra X1 256 autorally 1.506 45±0.575 889uncertain1.506450.5758891.506\,45$\pm$ 0.575\,889start_ARG 1.506 45 end_ARG ± start_ARG 0.575 889 end_ARG Jetson Nano Tegra X1 256 MPPI-Generic 1.383 92±0.061 362 6uncertain1.383920.06136261.383\,92$\pm$ 0.061\,362\,6start_ARG 1.383 92 end_ARG ± start_ARG 0.061 362 6 end_ARG
@@
Jetson Nano Tegra X1 512 autorally 1.846 51±0.574 576uncertain1.846510.5745761.846\,51$\pm$ 0.574\,576start_ARG 1.846 51 end_ARG ± start_ARG 0.574 576 end_ARG Jetson Nano Tegra X1 512 MPPI-Generic 2.122 54±0.119 91uncertain2.122540.119912.122\,54$\pm$ 0.119\,91start_ARG 2.122 54 end_ARG ± start_ARG 0.119 91 end_ARG
@@
Jetson Nano Tegra X1 1024 autorally 3.244 59±0.607 794uncertain3.244590.6077943.244\,59$\pm$ 0.607\,794start_ARG 3.244 59 end_ARG ± start_ARG 0.607 794 end_ARG Jetson Nano Tegra X1 1024 MPPI-Generic 4.105 41±1.202 12uncertain4.105411.202124.105\,41$\pm$ 1.202\,12start_ARG 4.105 41 end_ARG ± start_ARG 1.202 12 end_ARG
@@
Jetson Nano Tegra X1 2048 autorally 5.921±0.577 712uncertain5.9210.5777125.921$\pm$ 0.577\,712start_ARG 5.921 end_ARG ± start_ARG 0.577 712 end_ARG Jetson Nano Tegra X1 2048 MPPI-Generic 7.296 89±0.398 892uncertain7.296890.3988927.296\,89$\pm$ 0.398\,892start_ARG 7.296 89 end_ARG ± start_ARG 0.398 892 end_ARG
@@
Jetson Nano Tegra X1 4096 autorally 11.1589±0.581 031uncertain11.15890.58103111.1589$\pm$ 0.581\,031start_ARG 11.1589 end_ARG ± start_ARG 0.581 031 end_ARG Jetson Nano Tegra X1 4096 MPPI-Generic 14.1242±0.133 369uncertain14.12420.13336914.1242$\pm$ 0.133\,369start_ARG 14.1242 end_ARG ± start_ARG 0.133 369 end_ARG
@@
Jetson Nano Tegra X1 6144 autorally 16.5458±0.605 249uncertain16.54580.60524916.5458$\pm$ 0.605\,249start_ARG 16.5458 end_ARG ± start_ARG 0.605 249 end_ARG Jetson Nano Tegra X1 6144 MPPI-Generic 21.0746±0.229 213uncertain21.07460.22921321.0746$\pm$ 0.229\,213start_ARG 21.0746 end_ARG ± start_ARG 0.229 213 end_ARG
@@
Jetson Nano Tegra X1 8192 autorally 21.7672±0.628 532uncertain21.76720.62853221.7672$\pm$ 0.628\,532start_ARG 21.7672 end_ARG ± start_ARG 0.628 532 end_ARG Jetson Nano Tegra X1 8192 MPPI-Generic 28.0223±0.260 565uncertain28.02230.26056528.0223$\pm$ 0.260\,565start_ARG 28.0223 end_ARG ± start_ARG 0.260 565 end_ARG
@@
Jetson Nano Tegra X1 16384 autorally 44.1579±0.947 064uncertain44.15790.94706444.1579$\pm$ 0.947\,064start_ARG 44.1579 end_ARG ± start_ARG 0.947 064 end_ARG Jetson Nano Tegra X1 16384 MPPI-Generic 56.6497±0.350 869uncertain56.64970.35086956.6497$\pm$ 0.350\,869start_ARG 56.6497 end_ARG ± start_ARG 0.350 869 end_ARG
Jetson Orin Nano (8 GB) 128 autorally 0.946 763±0.292 363uncertain0.9467630.2923630.946\,763$\pm$ 0.292\,363start_ARG 0.946 763 end_ARG ± start_ARG 0.292 363 end_ARG Jetson Orin Nano (8 GB) 128 MPPI-Generic 0.629 243±0.1296uncertain0.6292430.12960.629\,243$\pm$ 0.1296start_ARG 0.629 243 end_ARG ± start_ARG 0.1296 end_ARG
Jetson Orin Nano (8 GB) 256 autorally 0.903 75±0.066 688 6uncertain0.903750.06668860.903\,75$\pm$ 0.066\,688\,6start_ARG 0.903 75 end_ARG ± start_ARG 0.066 688 6 end_ARG Jetson Orin Nano (8 GB) 256 MPPI-Generic 0.607 396±0.022 103 4uncertain0.6073960.02210340.607\,396$\pm$ 0.022\,103\,4start_ARG 0.607 396 end_ARG ± start_ARG 0.022 103 4 end_ARG
Jetson Orin Nano (8 GB) 512 autorally 0.909 514±0.033 216 3uncertain0.9095140.03321630.909\,514$\pm$ 0.033\,216\,3start_ARG 0.909 514 end_ARG ± start_ARG 0.033 216 3 end_ARG Jetson Orin Nano (8 GB) 512 MPPI-Generic 0.728 221±0.030 154 5uncertain0.7282210.03015450.728\,221$\pm$ 0.030\,154\,5start_ARG 0.728 221 end_ARG ± start_ARG 0.030 154 5 end_ARG
Jetson Orin Nano (8 GB) 1024 autorally 1.031 62±0.030 783 8uncertain1.031620.03078381.031\,62$\pm$ 0.030\,783\,8start_ARG 1.031 62 end_ARG ± start_ARG 0.030 783 8 end_ARG Jetson Orin Nano (8 GB) 1024 MPPI-Generic 1.030 02±0.034 969 4uncertain1.030020.03496941.030\,02$\pm$ 0.034\,969\,4start_ARG 1.030 02 end_ARG ± start_ARG 0.034 969 4 end_ARG
Jetson Orin Nano (8 GB) 2048 autorally 1.408 94±0.055 614 6uncertain1.408940.05561461.408\,94$\pm$ 0.055\,614\,6start_ARG 1.408 94 end_ARG ± start_ARG 0.055 614 6 end_ARG Jetson Orin Nano (8 GB) 2048 MPPI-Generic 1.645 41±0.108 883uncertain1.645410.1088831.645\,41$\pm$ 0.108\,883start_ARG 1.645 41 end_ARG ± start_ARG 0.108 883 end_ARG
Jetson Orin Nano (8 GB) 4096 autorally 2.712 73±0.124 459uncertain2.712730.1244592.712\,73$\pm$ 0.124\,459start_ARG 2.712 73 end_ARG ± start_ARG 0.124 459 end_ARG Jetson Orin Nano (8 GB) 4096 MPPI-Generic 3.092 44±0.079 106 5uncertain3.092440.07910653.092\,44$\pm$ 0.079\,106\,5start_ARG 3.092 44 end_ARG ± start_ARG 0.079 106 5 end_ARG
Jetson Orin Nano (8 GB) 6144 autorally 3.862 98±0.100 405uncertain3.862980.1004053.862\,98$\pm$ 0.100\,405start_ARG 3.862 98 end_ARG ± start_ARG 0.100 405 end_ARG Jetson Orin Nano (8 GB) 6144 MPPI-Generic 4.545 14±0.135 722uncertain4.545140.1357224.545\,14$\pm$ 0.135\,722start_ARG 4.545 14 end_ARG ± start_ARG 0.135 722 end_ARG
Jetson Orin Nano (8 GB) 8192 autorally 5.255 01±0.131 005uncertain5.255010.1310055.255\,01$\pm$ 0.131\,005start_ARG 5.255 01 end_ARG ± start_ARG 0.131 005 end_ARG Jetson Orin Nano (8 GB) 8192 MPPI-Generic 6.102 01±0.153 504uncertain6.102010.1535046.102\,01$\pm$ 0.153\,504start_ARG 6.102 01 end_ARG ± start_ARG 0.153 504 end_ARG
Jetson Orin Nano (8 GB) 16384 autorally 9.940 03±0.196 983uncertain9.940030.1969839.940\,03$\pm$ 0.196\,983start_ARG 9.940 03 end_ARG ± start_ARG 0.196 983 end_ARG Jetson Orin Nano (8 GB) 16384 MPPI-Generic 12.0695±0.217 526uncertain12.06950.21752612.0695$\pm$ 0.217\,526start_ARG 12.0695 end_ARG ± start_ARG 0.217 526 end_ARG